How do I solve 1-9/x=7 ?
Answer by checkley77(12569)

1-9/x=7
-9/x=7-1
-9/x=6
6x=-9
x=-9/6
x=-3/2 ans.
Proof:
1-9/(-3/2)=7
1-9*-2/3)=7
1+18/3=7
1+6=7
7=7